# Strand, London WC2R 2LS, UK

### v1

@@ 3.3.1 EMGB AdS Black Hole
Taking $\beta\rightarrow 0$, to leading order gives that the first term in the denominator dominates and thus to leading order, the Bekenstein-Hawking entropy is the same as in the zero-charge case, e.g., see Chamblin:1999tk . We will provide further remarks on this in section 5.
@@ 4 Lorentzian Linearized Dynamics
The IBVP, compared to the Euclidean case, is much more subtle. For instance, the (Euclidean) ellipticity of conformal boundary conditions in Einstein gravity was already proven in ref. anderson2008boundary , while results on the initial boundary corner value problem in Lorentzian signature are quite recent An:2021fcq ; Liu:2025xij ; An:2025cbs . Adding higher-curvature corrections can only make these proofs more involved. In fact, for EGB, the initial value problem has only recently been understood Kovacs:2020pns ; Kovacs:2020ywu ; Figueras:2024bba , [^19] This recent progress built on earlier work by Choquet-Bruhat Choquet-Bruhat:1988jdt ; choquet1989gravitation , who showed that the IVP of EGB gravity is well-posed in the restrictive case of analytic initial data. and we are not aware of any mathematical results regarding the problem in the presence of boundaries.
@@
The interaction between bulk and boundary degrees of freedom is encoded in the generalized (conformal) version of the Brown-York stress tensor, that in EGB is found to be [^20] In defining the conformal stress tensor we followed the conventions of ref. Anninos:2024xhc .
$T_{\mu\nu}^{(\text{EGB})}=-\dfrac{e^{(d-2){\boldsymbol{\omega}}}}{8\pi G_{N}}\left(\mathcal{K}^{\mu\nu}-\frac{1}{d}h^{\mu\nu}\mathcal{K}+\frac{\alpha}{(d-2)(d-3)}\Big(\mathcal{B}^{\mu\nu}-\dfrac{1}{d}h^{\mu\nu}\mathcal{B}\Big)\right)\ ,$ (78)
where $\mathcal{B}=h^{\mu\nu}\mathcal{B}_{\mu\nu}$, and the (long) explicit form of the tensor $\mathcal{B}_{\mu\nu}$ is given in eq. (117). For the Einstein case, this conformal stress tensor was derived in refs. An:2021fcq ; Odak:2021axr . For a derivation of the EGB generalization, see appendix A, e.g., eqs. (105) and (133).
@@
where the divergence is taken with respect to the conformal representative of the metric at the boundary $g]_{\mu\nu}$. In Einstein gravity, the stress tensor is transverse provided $\mathcal{K}$ is fixed to a constant at the boundary Anninos:2024xhc . In EGB, we observe that this generalizes and the conformal stress tensor is covariantly conserved when $\mathcal{P}$ is constant along $\partial\mathcal{M}$.
@@
The equivalent of the Hamiltonian constraint, for a boundary metric of the form (77), becomes a dynamical equation for the boundary mode ${\boldsymbol{\omega}}(y^{m})$. In Einstein gravity, this can be nicely written in terms of the conformal stress tensor, see e.g., eq. (2.13) of ref. Anninos:2024xhc . It would be interesting to see whether a similar formula can be obtained for EGB gravity. Given it is cumbersome, we will not attempt to write here the Hamiltonian constraint as a function of the ${\boldsymbol{\omega}}(y^{m})$ and $g]_{\mu\nu}$ in full generality.
@@
Instead, we will solve a simplified version of the problem Anninos:2024wpy ; Liu:2024ymn ; Anninos:2024xhc ; Galante:2025tnt . We will consider the bulk solution to be the Lorentzian version of the $f(r)$ metrics described in section 2, see eq. (17). Given the previous discussion, in the presence of a finite boundary subject to (extended) conformal boundary conditions, this does not fully solve the gravitational problem, as the Weyl factor at the boundary is still allowed to fluctuate. We will provide the form of the non-linear equation governing the dynamics of the Weyl factor when the conformal representative of the metric at the boundary is chosen to be the standard metric on $\mathbb{R}\times\Sigma_{k,d-1}$ (up to the overall Weyl factor which is only allowed to depend on the boundary time). Then we will consider the linearized version of that equation, expanded around the static solution with ${\boldsymbol{\omega}}=0$. For Einstein gravity with spherical boundaries, the solutions to the linearized problem are always exponentially increasing (and decreasing) with time Galante:2025tnt . In EGB, we show that this is still true for the spherical case. We also show that planar and hyperbolic boundaries allow for oscillatory behavior, even in Einstein gravity. We will consider both the charged and uncharged cases.
@@
Before moving into that, let us make a brief comment about the bulk dynamics problem. Of course, in the same way we are linearizing the boundary dynamics equation, we could consider the bulk linearized problem. Around spherically symmetric backgrounds in Einstein gravity, this problem has been analyzed in refs. Anninos:2023epi ; Anninos:2024wpy ; Liu:2024ymn ; Anninos:2024xhc . Namely, the linearized perturbations admit a Kodama-Ishibashi decomposition Kodama:2000fa ; Kodama:2003jz , which results in dynamics governed by a master field, whose equation of motion is solvable for maximally symmetric spacetimes. In a Fourier basis, imposing conformal boundary conditions selects the allowed mode frequencies. It was observed that this problem in Einstein gravity admits solutions which grow exponentially with time at the linearized level Anninos:2023epi ; Anninos:2024wpy ; Liu:2024ymn ; Anninos:2024xhc ; Liu:2025xij .
@@ 4.1 Boundary mode dynamics
In the limit $\alpha\rightarrow 0$ we recover the corresponding expression from Einstein gravity Liu:2024ymn ; Galante:2025tnt , namely eqs. (198) and (205) for planar and hyperbolic horizons, respectively. Note that a solution to eq. (82) leaves the bulk solution (80) unchanged, so these solutions can all be interpreted as physical diffeomorphims. Moreover, as a second order (non-linear) differential equation, eq. (82) needs to be supplemented by corner data, $\{{\boldsymbol{\omega}}(u)|_{\mathcal{C}},\partial_{u}{\boldsymbol{\omega}}(u)|_{\mathcal{C}}\}$. These data is neither part of the standard initial data of the initial value problem nor of the extended conformal boundary conditions and resides at what is usually termed the corner, i.e., the intersection between the Cauchy slice and the timelike boundary. After imposing this corner data, together with the standard initial data (at least within Einstein gravity), solutions to the initial boundary corner value problem (IBCVP) exist and are unique Anninos:2022ujl ; An:2025rlw . We expect that something similar holds for the IBCVP in EGB gravity, subject to the ECBC.
@@
We remark that interestingly, in pure Einstein-Maxwell theory subject to CBC (3), when the charge is large enough the behavior of $\lambda^{2}(\mathfrak{r})$ closely resembles that of the curves in the right plot of figure 5. This is demonstrated in appendix G. Thus, all the stable black brane modes discussed within EMGB exist also in EM, with the effect of the Gauss-Bonnet correction being that now stable modes start appearing when the boundary is placed further away from the horizon. In addition, the near-horizon growing modes can be associated to a Rindler-like behavior. In particular, taking the planar Rindler line element to be,
@@
for some constant $c$ and preforming a similar analysis subject to the extended conformal boundary conditions (81), one observes that the linearized dynamics of the corresponding conformal factor has frequencies
$\lambda=\pm\frac{\sqrt{d(\mathfrak{r}^{2}/\ell^{2}-2\tilde{\alpha})}}{\sqrt{(\mathfrak{r}^{2}/\ell^{2}+2\tilde{\alpha})}\ell}\$. (89)
For $0\leq\tilde{\alpha}<\tfrac{1}{4}$, it turns out that the radius of the outer horizon $r_{h}$ is such that $\lambda\in\mathbb{R}$ and only exponentially growing (and decaying) modes appear. The corresponding growing modes when $\tilde{\alpha}<0$ can also be linked to this behavior. The observation that exponentially growing modes appear to be a generic feature of near-horizon physics subject to CBC was already made in the context of Einstein gravity e.g., in ref. Anninos:2024wpy .
@@ 5 Discussion
In Einstein gravity, spherical cavities around maximally symmetric spacetimes allow for exponentially growing modes, that grow faster for large angular momentum  Anninos:2023epi ; Anninos:2024wpy ; Liu:2024ymn ; Anninos:2024xhc ; Liu:2025xij . These large angular momentum modes have radial profiles that localize exponentially close to the boundary. One might imagine at these scales higher curvature (or quantum) effects might affect the results obtained in Einstein gravity. In this work we studied the behavior of the spherically symmetric modes, but a systematic study of the large angular momentum modes is also possible. Allowed frequencies will be now obtained by imposing the generalized conformal boundary conditions for EGB. It would be interesting to see how a perturbative correction in the GB coupling might change the behavior found in Einstein gravity of the large angular momentum modes. The Kodama-Ishibashi formalism has been extended to LL gravity Ishibashi:2011ws , which might serve useful for this purpose.
@@
In ref. Liu:2024ymn , a one-parameter family of boundary conditions within Einstein gravity was introduced. In particular, the authors considered putting an arbitrary numerical coefficient $\Theta$ in front of the Gibbons-Hawking-York boundary term, i.e., $\frac{\Theta}{8\pi G_{N}}\int_{\partial\mathcal{M}}d^{d}y\sqrt{h}\mathcal{K}\ ,$ such that the boundary conditions fix the conformal class of the metric at the boundary, and some combination of the determinant of the induced metric and the trace of the extrinsic curvature, i.e., $\{[h_{mn}]|_{\partial\mathcal{M}},h^{p}\mathcal{K}|_{\partial\mathcal{M}}\}$, such that $p$ is related to $\Theta$ through $\Theta=1+\tfrac{d-1}{d(2p-1)}$. [^22] This modified action is also stationary under the following generalized umbilic boundary conditions $\{\tilde{\pi}^{\mu\nu}|_{\partial\mathcal{M}}=0,h^{p}\mathcal{K}|_{\partial\mathcal{M}}\}$. The $p=0$ case reduces to the conformal boundary conditions of Einstein gravity. One can generalize the extended boundary conditions discussed in this paper by allowing a general coefficient in front of the GB boundary term, $\frac{\Phi}{16\pi G_{N}}\int_{\partial\mathcal{M}}d^{d}y\sqrt{h}\mathcal{Q}$. This action has a well-defined variational principal provided the following quantities are kept fixed at the boundary, $\{[h_{mn}]|_{\partial\mathcal{M}},h^{p}\mathcal{P}|_{\partial\mathcal{M}}\}$, where $\Theta$ is chosen as above and $\Phi$ is given by $\Phi=1+\tfrac{d-3}{d(2p-1)}$. The detailed derivation is given in appendix A. It would be interesting to analyze the thermodynamics stemming from this generalized action and the corresponding generalized boundary conditions. Note that at high temperatures, the behavior will not be extensive unless $p=0$, following the argument of ref. Banihashemi:2024yye given for Einstein gravity.
@@
To obtain the results presented in this paper we had to make use of the natural generalization of the conformal boundary conditions (3) to Gauss-Bonnet gravity, namely the extended conformal boundary conditions (12). The difference is in the scalar quantity that is fixed at the boundary (along with the conformal class of the induced metric) is a higher derivative extension of the trace of the extrinsic curvature. Similarly, for more general Lanczos-Lovelock (LL) theories, this scalar quantity will receive further higher curvature contributions. In line with what was observed here for GB gravity, we expect these higher curvature corrections will be proportional to the boundary terms constructed in Myers:1987yn for the LL theory subject to Dirichlet boundary conditions. More specifically, recall that the action of the $2n$-th LL theory corresponds to the $2n$-dimensional Euler density. [^23] For example, the Einstein-Hilbert term is obtained for $n=1$, Gauss-Bonnet for $n=2$ and so on. The boundary term is then the contribution to the Euler density if the $2n$-dimensional manifold has a boundary.
@@
One may also consider more general higher–curvature interactions, e.g., the $R^{4}$ terms appearing in type II superstring theory Gross:1986iv ; Grisaru:1986vi ; Myers:1987qx . In general, however, such corrections do not preserve the property that the gravitational equations of motion remain second order in time derivatives. Consequently, the underlying classical system of non-linear partial differential equations requires an augmented initial value formulation in which higher (time) derivatives of the metric, or equivalently the associated higher conjugate momenta, must be specified Ostrogradsky:1850fid . An alternative viewpoint is that generic higher–curvature terms introduce additional propagating degrees of freedom (typically ghosts) at the scale set by their couplings. In the effective field theory approach, one does not wish to retain these modes but only to capture how the long-wavelength ‘Einsteinian’ modes are perturbed by the higher–curvature interactions. Projecting out the unwanted excitations demands a more refined treatment, e.g., see Simon:1990ic ; Parker:1993dk , but would not substantially alter the qualitative structure of the boundary data. The special virtue of the EGB and, more generally, Lovelock theories is that the gravitational equations of motion remain second order in time derivatives. Accordingly, constructing the extended conformal boundary conditions (12) was comparatively straightforward. For generic higher–curvature corrections treated perturbatively around Einstein gravity, we expect that an analogous set of ECBC can be defined in which, as in eq. (13), the boundary data is the conformal class of the boundary metric and a scalar built from the extrinsic curvature augmented by the appropriate higher–curvature corrections.
@@
Additionally, from an effective field theory point of view, it is natural to consider higher derivative terms involving the gauge field (and possibly mixed terms combining the gauge field strength and the curvature tensor). Black hole solutions with four derivative terms in the Maxwell field were studied in GB gravity Wiltshire:1988uq . Another natural extension of the present work might be the dimensional reduction of GB gravity, which in the lower dimensional theory, extends the GB term is a special family of higher derivative terms involving the Kaluza-Klein gauge field while preserving the second-order character of the equations of motion, e.g., see Mueller-Hoissen:1985www ; Huang:1988mw . It would be interesting to consider conformal thermodynamics for black holes in this theory.
@@
We would also like to comment here on the following: [^24] We thank D. Anninos for illuminating discussions on this point. the gravitational path integral must, in principle, be invariant under field redefinitions. Under such transformations, however, the effective couplings appearing in the action (9) are generally reshuffled. It is therefore necessary to specify what one means, in a covariant sense, by a particular coupling such as the Gauss–Bonnet parameter $\alpha$. The natural prescription is to identify its distinct contribution to an appropriate physical observable, e.g., a boundary graviton correlator. For instance, in the analysis of EGB gravity within the AdS/CFT framework Buchel:2009sk , the Gauss–Bonnet coupling controls a unique tensor structure in the three-point function of the boundary stress tensor. A context in which such a covariant definition of higher–curvature couplings plays an essential role is the computation of the de Sitter horizon entropy using the sphere path integral Anninos:2020hfj .
@@
Note the overall scaling behavior of the flat space conformal entropy (67) can be understood through dimensional analysis. In particular, in the case of Einstein gravity, the only possible dimensionless combination is $G_{N}\mathcal{K}^{d-1}$ and therefore it seems that all the coefficients of the thermal effective action (29) will have the form $c_{i}(d)/G_{N}\mathcal{K}^{d-1}$, with some dimension-dependent numerical coefficient $c_{i}(d)$. That is true for the first three coefficients, as computed in Banihashemi:2025qqi , and our argument suggests that it is true for every order in the small-$\tilde{\beta}$ expansion.
@@
The extensivity of the high-temperature conformal thermodynamics subject to (E)CBC is hinting toward a dual field theoretic description that is local, in contrast to the $T\bar{T}$-deformed field theories dual to finite Dirichlet boundaries in AdS/CFT duality Freidel:2008sh ; McGough:2016lol ; Zamolodchikov:2004ce ; Smirnov:2016lqw ; Hartman:2018tkw . Moreover, the extensive behavior is universal and does not depend on the sign of the cosmological constant. This curious observation might serve as a guiding principle towards finding microscopic realizations of the conformal boundaries. For example, ref. Allameh:2025gsa recently proposed a dual theory for $d=2$ with negative cosmological constant. To gain intuition for the higher dimensional case, one could attempt to compute stress tensor and/or matter correlation functions inside the worldtube and interpret them in terms of boundary data. Note that, as discussed in Anninos:2024xhc , the presence of the scale $\mathcal{P}$ might induce some sort of conformal symmetry breaking, at least when the boundary is sufficiently close to the conformal boundary of AdS.
In addition, it would be interesting to study whether the familiar notions of causal Hamilton:2005ju and entanglement Dong:2016eik wedge reconstruction from AdS/CFT generalize to conformal boundaries placed at finite distance. This could be achieved by leveraging the proposal for generalized entanglement wedges for gravitating spacetime regions, that was put forward in Bousso:2022hlz ; Bousso:2023sya . Ultimately, one would wish to understand quantum aspects of conformal timelike boundaries and their role within string/M-theory. Liouville walls in the string target space discussed in e.g., Silverstein:2022dfj , seem to be an intriguing starting point for progress in this direction.

### v2

@@ 3.3.1 EMGB AdS Black Hole
Taking $\beta\rightarrow 0$, to leading order gives that the first term in the denominator dominates and thus to leading order, the Bekenstein-Hawking entropy is the same as in the zero-charge case, e.g., see Chamblin et al. (1999). We will provide further remarks on this in section 5.
@@ 4 Lorentzian Linearized Dynamics
The IBVP, compared to the Euclidean case, is much more subtle. For instance, the (Euclidean) ellipticity of conformal boundary conditions in Einstein gravity was already proven in ref. Anderson (2008), while results on the initial boundary corner value problem in Lorentzian signature are quite recent An and Anderson (2021); Liu et al. (2025); An and Anderson (2025c). Adding higher-curvature corrections can only make these proofs more involved. In fact, for EGB, the initial value problem has only recently been understood Kovács and Reall (2020b, a); Figueras et al. (2024), [^20] This recent progress built on earlier work by Choquet-Bruhat Choquet-Bruhat (1988, 1989), who showed that the IVP of EGB gravity is well-posed in the restrictive case of analytic initial data. and we are not aware of any mathematical results regarding the problem in the presence of boundaries.
@@
The interaction between bulk and boundary degrees of freedom is encoded in the generalized (conformal) version of the Brown-York stress tensor, that in EGB is found to be [^21] In defining the conformal stress tensor we followed the conventions of ref. Anninos et al. (2025a).
$T_{\mu\nu}^{(\text{EGB})}=-\dfrac{e^{(d-2){\boldsymbol{\omega}}}}{8\pi G_{N}}\left(\mathcal{K}_{\mu\nu}-\frac{1}{d}h_{\mu\nu}\mathcal{K}+\frac{\alpha}{(d-2)(d-3)}\Big(\mathcal{B}_{\mu\nu}-\dfrac{1}{d}h_{\mu\nu}\mathcal{B}\Big)\right)\ ,$ (78)
where $\mathcal{B}=h^{\mu\nu}\mathcal{B}_{\mu\nu}$, and the (long) explicit form of the tensor $\mathcal{B}_{\mu\nu}$ is given in eq. (116). For the Einstein case, this conformal stress tensor was derived in refs. An and Anderson (2021); Odak and Speziale (2021). For a derivation of the EGB generalization, see appendix A, e.g., eqs. (105) and (132).
@@
where the divergence is taken with respect to the conformal representative of the metric at the boundary $g]_{\mu\nu}$. In Einstein gravity, the stress tensor is transverse provided $\mathcal{K}$ is fixed to a constant at the boundary Anninos et al. (2025a). In EGB, we observe that this generalizes and the conformal stress tensor is covariantly conserved when $\mathcal{P}$ is constant along $\partial\mathcal{M}$.
@@
The equivalent of the Hamiltonian constraint, for a boundary metric of the form (77), becomes a dynamical equation for the boundary mode ${\boldsymbol{\omega}}(y^{m})$. In Einstein gravity, this can be nicely written in terms of the conformal stress tensor, see e.g., eq. (2.13) of ref. Anninos et al. (2025a). It would be interesting to see whether a similar formula can be obtained for EGB gravity. Given it is cumbersome, we will not attempt to write here the Hamiltonian constraint as a function of the ${\boldsymbol{\omega}}(y^{m})$ and $g]_{\mu\nu}$ in full generality.
@@
Instead, we will solve a simplified version of the problem Anninos et al. (2024); Liu et al. (2024); Anninos et al. (2025a); Galante et al. (2025). We will consider the bulk solution to be the Lorentzian version of the $f(r)$ metrics described in section 2, see eq. (17). Given the previous discussion, in the presence of a finite boundary subject to (extended) conformal boundary conditions, this does not fully solve the gravitational problem, as the Weyl factor at the boundary is still allowed to fluctuate. We will provide the form of the non-linear equation governing the dynamics of the Weyl factor when the conformal representative of the metric at the boundary is chosen to be the standard metric on $\mathbb{R}\times\Sigma_{k,d-1}$ (up to the overall Weyl factor which is only allowed to depend on the boundary time). Then we will consider the linearized version of that equation, expanded around the static solution with ${\boldsymbol{\omega}}=0$. For Einstein gravity with spherical boundaries, the solutions to the linearized problem are always exponentially increasing (and decreasing) with time Galante et al. (2025). In EGB, we show that this is still true for the spherical case. We also show that planar and hyperbolic boundaries allow for oscillatory behavior, even in Einstein gravity. We will consider both the charged and uncharged cases.
@@
Before moving into that, let us make a brief comment about the bulk dynamics problem. Of course, in the same way we are linearizing the boundary dynamics equation, we could consider the bulk linearized problem. Around spherically symmetric backgrounds in Einstein gravity, this problem has been analyzed in refs. Anninos et al. (2023a, 2024); Liu et al. (2024); Anninos et al. (2025a). Namely, the linearized perturbations admit a Kodama-Ishibashi decomposition Kodama et al. (2000); Kodama and Ishibashi (2003), which results in dynamics governed by a master field, whose equation of motion is solvable for maximally symmetric spacetimes. In a Fourier basis, imposing conformal boundary conditions selects the allowed mode frequencies. It was observed that this problem in Einstein gravity admits solutions which grow exponentially with time at the linearized level Anninos et al. (2023a, 2024); Liu et al. (2024); Anninos et al. (2025a); Liu et al. (2025).
@@ 4.1 Boundary mode dynamics
In the limit $\alpha\rightarrow 0$ we recover the corresponding expression from Einstein gravity Liu et al. (2024); Galante et al. (2025), namely eqs. (197) and (204) for planar and hyperbolic horizons, respectively. Note that a solution to eq. (82) leaves the bulk solution (80) unchanged, so these solutions can all be interpreted as physical diffeomorphims. Moreover, as a second order (non-linear) differential equation, eq. (82) needs to be supplemented by corner data, $\{{\boldsymbol{\omega}}(u)|_{\mathcal{C}},\partial_{u}{\boldsymbol{\omega}}(u)|_{\mathcal{C}}\}$. These data is neither part of the standard initial data of the initial value problem nor of the extended conformal boundary conditions and resides at what is usually termed the corner, i.e., the intersection between the Cauchy slice and the timelike boundary. After imposing this corner data, together with the standard initial data (at least within Einstein gravity), solutions to the initial boundary corner value problem (IBCVP) exist and are unique Anninos et al. (2023b); An and Anderson (2025a). We expect that something similar holds for the IBCVP in EGB gravity, subject to the ECBC.
@@
We remark that interestingly, in pure Einstein-Maxwell theory subject to CBC (3), when the charge is large enough the behavior of $\lambda^{2}(\mathfrak{r})$ closely resembles that of the curves in the right plot of figure 5. This is demonstrated in appendix G. Thus, all the stable black brane modes discussed within EMGB exist also in EM, with the effect of the Gauss-Bonnet correction being that now stable modes start appearing when the boundary is placed further away from the horizon. In addition, the near-horizon growing modes can be associated to a Rindler-like behavior Anninos et al. (2025b). In particular, taking the planar Rindler line element to be,
@@
for some constant $c$ and performing a similar analysis subject to the extended conformal boundary conditions (81), one observes that the linearized dynamics of the corresponding conformal factor has frequencies
$\lambda\ell=\pm 1+\mathcal{O}\Big(\frac{\tilde{\alpha}}{\mathfrak{r}/\ell^{2}}\Big)\ ,$ (89)
which is the Einstein gravity behavior, and higher-derivative effects kick in once we move slightly away from the strict Rindler limit. Consequently, all the near-horizon growing modes observed in the context of EMGB can be traced back to such a behavior. We further remark that the expansion in eq. (89), including all the sub-leading terms, mimics structurally the (Rindler-like) stretched horizon limit explored in ref. Anninos et al. (2025b), see e.g., eq. (5.2) there. The observation that exponentially growing modes appear to be a generic feature of near-horizon physics subject to CBC was already made in the context of Einstein gravity, e.g., in ref. Anninos et al. (2024).
@@ 5 Discussion
In Einstein gravity, spherical cavities around maximally symmetric spacetimes allow for exponentially growing modes, that grow faster for large angular momentum  Anninos et al. (2023a, 2024); Liu et al. (2024); Anninos et al. (2025a); Liu et al. (2025); Anninos et al. (2025b). These large angular momentum modes have radial profiles that localize exponentially close to the boundary. One might imagine at these scales higher curvature (or quantum) effects might affect the results obtained in Einstein gravity. In this work we studied the behavior of the spherically symmetric modes, but a systematic study of the large angular momentum modes is also possible. Allowed frequencies will be now obtained by imposing the generalized conformal boundary conditions for EGB (12). It would be interesting to see how the corrections in the GB coupling affect the behavior of the large angular momentum modes, found in Einstein gravity. The Kodama-Ishibashi formalism has been extended to LL gravity Ishibashi and Kodama (2011), which might serve useful for this purpose.
@@
In ref. Liu et al. (2024), a one-parameter family of boundary conditions within Einstein gravity was introduced. In particular, the authors considered putting an arbitrary numerical coefficient $\Theta$ in front of the Gibbons-Hawking-York boundary term, i.e., $\frac{\Theta}{8\pi G_{N}}\int_{\partial\mathcal{M}}d^{d}y\sqrt{h}\mathcal{K}\ ,$ such that the boundary conditions fix the conformal class of the metric at the boundary, and some combination of the determinant of the induced metric and the trace of the extrinsic curvature, i.e., $\{[h_{mn}]|_{\partial\mathcal{M}},h^{p}\mathcal{K}|_{\partial\mathcal{M}}\}$, such that $p$ is related to $\Theta$ through $\Theta=1+\tfrac{d-1}{d(2p-1)}$. [^23] This modified action is also stationary under the following generalized umbilic boundary conditions $\{\tilde{\pi}^{\mu\nu}|_{\partial\mathcal{M}}=0,h^{p}\mathcal{K}|_{\partial\mathcal{M}}=\text{fixed}\}$ . The $p=0$ case reduces to the conformal boundary conditions of Einstein gravity. One can generalize the extended boundary conditions discussed in this paper by allowing a general coefficient in front of the GB boundary term, $\frac{\Phi}{16\pi G_{N}}\int_{\partial\mathcal{M}}d^{d}y\sqrt{h}\mathcal{Q}$. This action has a well-defined variational principal provided the following quantities are kept fixed at the boundary, $\{[h_{mn}]|_{\partial\mathcal{M}},h^{p}\mathcal{P}|_{\partial\mathcal{M}}\}$, where $\Theta$ is chosen as above and $\Phi$ is given by $\Phi=1+\tfrac{d-3}{d(2p-1)}$. The detailed derivation is given in appendix A. It would be interesting to analyze the thermodynamics stemming from this generalized action and the corresponding generalized boundary conditions. Note that at high temperatures, the behavior will not be extensive unless $p=0$, following the argument of ref. Banihashemi et al. (2025a) given in the context of Einstein gravity.
@@
To obtain the results presented in this paper we had to make use of the natural generalization of the conformal boundary conditions (3) to Gauss-Bonnet gravity, namely the extended conformal boundary conditions (12). The difference is in the scalar quantity that is fixed at the boundary (along with the conformal class of the induced metric) is a higher derivative extension of the trace of the extrinsic curvature. Similarly, for more general Lanczos-Lovelock theories, this scalar quantity will receive further higher curvature contributions. In line with what was observed here for GB gravity, we expect these higher curvature corrections will be proportional to the boundary terms constructed in Myers (1987a) for the LL theory subject to Dirichlet boundary conditions. More specifically, recall that the action of the $2n$-th LL theory corresponds to the $2n$-dimensional Euler density. [^24] For example, the Einstein-Hilbert term is obtained for $n=1$, Gauss-Bonnet for $n=2$ and so on. The boundary term is then the contribution to the Euler density if the $2n$-dimensional manifold has a boundary.
@@
One may also consider more general higher–curvature interactions, e.g., the $R^{4}$ terms appearing in type II superstring theory Gross and Witten (1986); Grisaru and Zanon (1986); Myers (1987b). In general, however, such corrections do not preserve the property that the gravitational equations of motion remain second order in time derivatives. Consequently, the underlying classical system of non-linear partial differential equations requires an augmented initial value formulation in which higher (time) derivatives of the metric, or equivalently the associated higher conjugate momenta, must be specified Ostrogradsky (1850). An alternative viewpoint is that generic higher–curvature terms introduce additional propagating degrees of freedom (typically ghosts) at the scale set by their couplings. In effective field theory, one does not wish to retain these modes but only to capture how the long-wavelength ‘Einsteinian’ modes are perturbed by the higher–curvature interactions. Projecting out the unwanted excitations demands a more refined treatment, e.g., see Simon (1990); Parker and Simon (1993), but would not substantially alter the qualitative structure of the boundary data. The special virtue of EGB and, more generally, Lovelock theories is that the gravitational equations of motion remain second order in time derivatives. Accordingly, constructing the extended conformal boundary conditions (12) was comparatively straightforward. For generic higher–curvature corrections treated perturbatively around Einstein gravity, we expect that an analogous set of ECBC can be defined in which, as in eq. (13), the boundary data is the conformal class of the boundary metric and a scalar built from the extrinsic curvature augmented by the appropriate higher–curvature corrections.
@@
Additionally, from an effective field theory point of view, it is natural to consider higher derivative terms involving the gauge field (and possibly mixed terms combining the gauge field strength and the curvature tensor). Black hole solutions with four derivative terms in the Maxwell field were studied in GB gravity Wiltshire (1988). Another natural extension of the present work might be the dimensional reduction of GB gravity, which in the lower dimensional theory, extends the GB term is a special family of higher derivative terms involving the Kaluza-Klein gauge field while preserving the second-order character of the equations of motion, e.g., see Mueller-Hoissen (1986); Huang (1988). It would be interesting to consider conformal thermodynamics for black holes in these theories, though we expect that the general high-temperature extensivity (4) persists.
@@
We would also like to comment here on the following: [^25] We thank D. Anninos for illuminating discussions on this point. the gravitational path integral must, in principle, be invariant under field redefinitions. Under such transformations, however, the effective couplings appearing in the action (9) are generally reshuffled. It is therefore necessary to specify what one means, in a covariant sense, by a particular coupling such as the Gauss–Bonnet parameter $\alpha$. The natural prescription is to identify its distinct contribution to an appropriate physical observable, e.g., a boundary graviton correlator. For instance, in the analysis of EGB gravity within the AdS/CFT framework Buchel et al. (2010), the Gauss–Bonnet coupling controls a unique tensor structure in the three-point function of the boundary stress tensor. A context in which such a covariant definition of higher–curvature couplings plays an essential role is the computation of the sphere path integral Anninos et al. (2022).
@@
Note the overall scaling behavior of the flat space conformal entropy (67) can be understood through dimensional analysis. In particular, in the case of Einstein gravity, the only possible dimensionless combination is $G_{N}\mathcal{K}^{d-1}$ and therefore it seems that all the coefficients of the thermal effective action (29) will have the form $c_{i}(d)/G_{N}\mathcal{K}^{d-1}$, with some dimension-dependent numerical coefficient $c_{i}(d)$. That is true for the first three coefficients, as computed in Banihashemi et al. (2025b), and our argument suggests that it is true for every order in the small-$\tilde{\beta}$ expansion.
@@
The extensivity of the high-temperature conformal thermodynamics subject to (E)CBC is hinting toward a dual field theoretic description that is local, in contrast to the $T\bar{T}$-deformed field theories dual to finite Dirichlet boundaries in AdS/CFT duality Freidel (2008); McGough et al. (2018); Zamolodchikov (2004); Smirnov and Zamolodchikov (2017); Hartman et al. (2019). Moreover, the extensive behavior is universal and does not depend on the sign of the cosmological constant. This curious observation might serve as a guiding principle towards finding microscopic realizations of these conformal boundaries. For example, ref. Allameh and Shaghoulian (2025b) recently proposed a dual theory for $d=2$ with negative cosmological constant. To gain intuition for the higher dimensional case, one could attempt to compute stress tensor and/or matter correlation functions inside the worldtube and interpret them in terms of boundary data. Note that, as discussed in Anninos et al. (2025a), the presence of the scale $\mathcal{P}$ might induce some sort of conformal symmetry breaking, at least when the boundary is sufficiently close to the conformal boundary of AdS.
In addition, it would be interesting to study whether the familiar notions of causal Hamilton et al. (2006) and entanglement Dong et al. (2016) wedge reconstruction from AdS/CFT generalize to conformal boundaries placed at finite distance. This could be achieved by leveraging the proposal for generalized entanglement wedges for gravitating spacetime regions, that was put forward in Bousso and Penington (2023a, b). Ultimately, one would wish to understand quantum aspects of conformal timelike boundaries and their role within string/M-theory. Liouville walls in the string target space discussed in e.g., Silverstein (2023), seem to be an intriguing starting point for progress in this direction.